i have a test tomorrow and im strugling can u help solve this question on my review
7n+2[3(l-n)-2(1+n)=14
another question is -10+4(3p+10)=18
one more is 66=-6/5(s+3)

Answer by stanbon(75887)

7n+2[3(l-n)-2(1+n)=14
You opened a bracket but you did not close it.
Best guess is:
7n+2[3-3n-2-2n]=14
7n+2[1-5n]=14
7n+2-10n=14
-3n=12
n=-4
-----------
another question is -10+4(3p+10)=18
-10+12p+40=18
12p+30=18
12p=-12
p=-1
---------------
one more is 66=-6/5(s+3)
66=-(6/5)(s+3)
330=-6(s+3)
-55=s+3
s=-58
---------
Cheers,
Stan H.